In a survey of 3270 adults aged 57 through 85 years, it was found that 86 9% of them used at least one prescription medication Complete parts (a) through (c) below.


a. How many of the 3270 subjects used at least one prescription medication?


Round to three decimal places.

make a conclusion
Answer #1

Given N = number of adults aged 57 through 85 years = 3270

Let p = percentage of adults aged 57 through 85 used at least one prescription medication.

p = 86.9% = 0.869

So , Number of adults used at least one prescription medication = N*p = 3270*0.869 = 2841.630

Number of adults used at least one prescription medication is 2842